# MASS: Motion-Aware Spatial–temporal Grounding for Physics Reasoning and Comprehension in Vision-Language Models

### v1

@@ A.3 Question Type Composition
Figure 5: Prompt template used for automatic evaluation of model answers against ground-truth references. The template presents the question, ground truth, and model output provided for LLM-as-a-judge evaluation and instructs the evaluator to output one of three outcomes—Correct, Incorrect, or Unclear—ensuring reliable and consistent scoring across predictions.
@@ Appendix E Experiments on Real-world Video QA
As a supplement to the experiments in the main paper, we provide additional results on two real-world video question-answering benchmarks, MMVU [zhao2025mmvu] and MVBench [li2024mvbench], to further assess the generalization ability of VLMs augmented with MASS. MMVU evaluates models on expert-level temporal, procedural, and interaction-centric reasoning, while MVBench is a widely used benchmark for temporal video understanding and action-centric tasks. Both datasets contain diverse video-based QA pairs captured from real-world scenarios.
Table 4 reports the performance of VLMs integrated with MASS alongside state-of-the-art baselines. On MMVU, large-scale close-source models achieve substantially higher accuracy due to the benchmark’s emphasis on expert-level, interdisciplinary reasoning—highlighting the limitations of smaller 7B models. Nevertheless, VLMs enhanced with MASS attain performance comparable to open-source models of similar scale, demonstrating strong generalization across tasks. On MVBench, models equipped with MASS consistently outperform all baselines, underscoring the effectiveness of motion-aware spatial–temporal grounding for general-purpose video understanding when capturing temporal information and leveraging such information in reasoning and question-answering.
@@
Qwen3-VL-8B 64.00 51.30
LLaVA-OneVision 50.41 55.24
@@
Table 4: Overall accuracies (%) on MMVU and MVBench. We report overall performance on two comprehensive, real-world video question-answering benchmarks, MMVU [zhao2025mmvu] and MVBench [li2024mvbench]. GPT-4o achieves the highest accuracy on both datasets. Notably, open-source VLMs augmented with MASS exhibit strong generalization, narrowing the gap to close-source models and transferring effectively from physics-oriented spatial–temporal grounding to broader video QA tasks.
Appendix F Case Study

### v2

@@ A.4 Motion Grounding Annotations
Figure 5: Prompt template used for automatic evaluation of model answers against ground-truth references. The template presents the question, ground truth, and model output provided for LLM-as-a-judge evaluation and guides the evaluator to produce one of three outcomes, Correct, Incorrect, or Unclear, ensuring reliable and consistent scoring across all predictions.
@@ Appendix E Experiments on Real-world Video QA
As a supplement to the experiments in the main paper, we provide additional results on two real-world video question-answering benchmarks, MMVU [67] and MVBench [30], to further assess the generalization ability of VLMs augmented with MASS. MMVU evaluates models on expert-level temporal, procedural, and interaction-centric reasoning, while MVBench is a widely used benchmark for temporal video understanding and action-centric tasks. Both datasets contain diverse video-based QA pairs captured from real-world scenarios.
Table 6 reports the performance of VLMs integrated with MASS alongside state-of-the-art baselines. On MMVU, large-scale close-source models achieve substantially higher accuracy due to the benchmark’s emphasis on expert-level, interdisciplinary reasoning—highlighting the limitations of smaller 7B models. Qwen2.5-VL-7B+MASS improves over its baseline on both benchmarks and narrows the gap to close-source models, while LLaVA-OneVision-7B+MASS shows a performance drop on these out-of-distribution tasks. As also shown in Table 8, LLaVA is more sensitive to detection and tracking quality: when salient entities are missed or motion cues are noisy—common in MMVU and MVBench—performance can degrade, motivating more selective motion-profile extraction in future work.
@@
Qwen2.5-VL-7B 64.00 51.30
LLaVA-OneVision-7B 50.41 55.24
@@
Table 6: Overall accuracies (%) on MMVU and MVBench. We report overall performance on two real-world video question-answering benchmarks, MMVU [67] and MVBench [30]. GPT-4o achieves the highest accuracy on both datasets. Qwen2.5-VL-7B+MASS improves over its baseline and narrows the gap to close-source models, while LLaVA-OneVision-7B+MASS is more sensitive to detection quality on these out-of-distribution tasks.
Appendix F Noise-Induced Robustness
We inject two noise types into the motion profile to test spatial–motion grounding robustness: (1) BBox/position noise, perturbing bounding boxes and initial/final 3D positions (SAM-2 & Grounding-DINO); (2) Trajectory Noise, perturbing 3D motion vectors (Co-Tracker). We apply a relatively large 10% Gaussian perturbation to the motion vectors, positions, and bounding boxes, both separately and jointly. Table 7 shows slight accuracy drops, yet MASS-augmented backbones still substantially outperform their baselines (Table 2 in the main paper) across all noise settings, even after RFT. Because our physical comprehension is mostly qualitative and the motion profiles serve as supplementary cues rather than exact measurements, the model tolerates moderate value-level perturbations. Failures mainly occur when noise flips the sign of motion components, i.e., reversing positive/negative values, leading the model to make wrong assumptions about motion directions and further cascading into a misinterpretation of the entire physical process.
Baselines (MASS) Noise Injection
Model (+ MASS) w/ w/o BBox Motion All
Qwen2.5-VL-7B 61.93 52.30 58.64 58.51 58.42
LLaVA-OneVision-7B 59.24 45.79 54.50 54.19 54.19
Table 7: Noise-induced robustness evaluation on MASS-augmented models. BBox, Motion, and All columns report accuracy under 10% Gaussian perturbations applied to bounding-box/position, 3D trajectory, and both components jointly. Baseline columns (w/ and w/o MASS) are provided as reference.
Appendix G Inference Time and Memory Usage
We report wall-clock inference time on 100 randomly sampled test cases on an NVIDIA L40S (Table 8). MASS increases per-query latency due to motion-profile extraction, while GPU memory increases only modestly. Our goal is to evaluate the benefit of explicit spatial–motion cues for physical understanding, not an optimized pipeline. Importantly, the motion profile is modular and cacheable per video, so it can be reused across multiple questions without re-extraction. Overhead can be further reduced via faster upstream modules and standard efficiency knobs (e.g., lower frame stride/resolution, smaller top-$K$ grounded entities), which we leave to future work.
Model Time (s/case) GPU Mem (GB)
w/o w/ w/o w/
LLaVA-OV-7B 1.86 6.49 14.97 15.69
Qwen2.5-VL-7B 5.83 10.46 15.45 16.18
Table 8: Per-query inference time (s) and GPU memory usage (GB) on an NVIDIA L40S over 100 randomly sampled test cases, with and without MASS.
Appendix H Multi-object Tracking and Camera Motion
We agree that dense tracking in crowded scenes increases compute and may introduce failures. MASS-Bench is entity-centric: we track only grounded target entities (top-$K$, tunable; default 5), and our QA design and video filtering avoid highly crowded, many-entity cases. Camera ego-motion remains an open challenge; the current benchmark largely contains limited camera motion to focus on entity-centric spatial/motion grounding. MASS does not assume perfect tracks and degrades gracefully under noisy motion cues (as shown in the noise robustness analysis in Table 7). We plan to extend the benchmark and method to crowded multi-entity tracking and stronger camera motion handling in future work.
Appendix I Case Study